OpenStudy (anonymous):

write an equation of the line containing the given point and perpendicular to the given line. (0,4); 9x+8y=5 The equation of the line is Y=____

OpenStudy (anonymous):

8y+9x=5 so y=(-9/8)x+5/8 so make the line y=(8/9)x+b through(0.4) so b=4 so the line: y=(8/9)x+4